OpenStudy (moneybaby980):

help What is the first step needed to solve 4 over 7 multiplied by x minus 5 equals negative 13? Subtract 13 from both sides Divide both sides by 7 Add 5 to both sides Multiply both sides by 4

OpenStudy (dtan5457):

\[\frac{ 4 }{ 7 }x-5=-13\] Whenever you want to solve for x..you want to isolate it, so that -5 should go to the other side.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

a) add \(5\) to both sides then b) multiply both sides by \(\frac{7}{4}\)